OpenStudy (anonymous):

Suppose we don't have a formula for g(x) but we know that g(2)=-4 and g'(x)+sqrt(x^2+5) for all x. (a) use linear approximation to estimate g(1.95) and g(2.05) (b) Are your estimates in part (a) too large or too small? Explain.

OpenStudy (anonymous):

my estimates were \[g(1.95)\approx-5.5 \]\[g(2.05)\approx-2.5\]

OpenStudy (anonymous):

i just don't get part b. if i had the graph i would look at the concavity but i dont

OpenStudy (kc_kennylau):

@hartnn

OpenStudy (anonymous):

Is g'(x)+sqrt(x^2+5)=0 ?

OpenStudy (anonymous):

g'(x) a function. g'(x) is the derivative of g(x)

OpenStudy (anonymous):

sorry it was a typo. g'(x)=sqrt(x^2+5)

OpenStudy (anonymous):

take out value of g from g(2)=-4 which is -8. replace -8 in the other equation for g

OpenStudy (anonymous):

where did you get -8

OpenStudy (anonymous):

you said you know that g(2)=-4, didnt you? solve it and you get -8. replace that for g in other equations.

OpenStudy (anonymous):

@muzzammil.raza

OpenStudy (phi):

for (b), look at the sign of g''(x) (the higher order terms of a taylor expansion)

OpenStudy (anonymous):

sorry websight failed.

OpenStudy (anonymous):

@phi i dont know what taylor expansion is. @ichiro what are you solving

OpenStudy (anonymous):

\[g'\left( x \right)=\int\limits \sqrt{x ^{2}+5}*1dx\] \[g \left( x \right)=\sqrt{x ^{2}+5}*x-\int\limits \frac{ 2x }{2\sqrt{x ^{2}+5} }*x dx\] \[g \left( x \right)=x \sqrt{x ^{2}+5}-\int\limits \frac{ x ^{2}+5-5 }{\sqrt{x ^{2}+5} }dx\]

OpenStudy (anonymous):

what is that curvy line thing

OpenStudy (anonymous):

\[g \left( x \right)=x \sqrt{x ^{2}+5}-\int\limits \sqrt{x ^{2}+5}dx+5\int\limits \frac{ dx }{\sqrt{x ^{2}+5} }\] \[g \left( x \right)=x \sqrt{x ^{2}+5}-g \left( x \right)+5 \int\limits \frac{ dx }{\sqrt{x ^{2}+5} }\]

OpenStudy (anonymous):

\[2g \left( x \right)=x \sqrt{x ^{2}+5}+5 I\]

OpenStudy (anonymous):

i don't understand any of this

OpenStudy (anonymous):

\[I=\int\limits \frac{ dx}{ \sqrt{x ^{2}+5} }\] Let me complete first.

OpenStudy (anonymous):

\[Putx=\sqrt{5}\tan \theta \] \[dx=\sqrt{5}\sec ^{2}\theta d \theta \]

OpenStudy (anonymous):

\[I=\int\limits \frac{ \sqrt{5}\sec ^{2}\theta d \theta }{\sqrt{5\tan ^{2}\theta+5} }\] \[I=\int\limits \frac{ \sqrt{5}\sec ^{2}\theta d \theta }{\sqrt{5}\sec \theta }=\int\limits \sec \theta d \theta \] \[=\ln \left| \sec \theta+\tan \theta \right|+c\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[I=\ln \left| \frac{ \sqrt{x ^{2}+5} }{ \sqrt{5} }+\frac{ x }{\sqrt{5} } \right|+c\]

OpenStudy (anonymous):

\[2g \left( x \right)=x \sqrt{x ^{2}+5}+5[\ln \left| \frac{ x+\sqrt{x ^{2}+5} }{\sqrt{5} } \right|+c\]

OpenStudy (anonymous):

\[g \left( x \right)=\frac{ x \sqrt{x ^{2}+5} }{ 2 }+\frac{ 5 }{ 2 }\ln \left| \frac{ x+\sqrt{x ^{2}+5} }{\sqrt{5} } \right|+\frac{ c }{ 2 }\] put x=2 and find c

OpenStudy (anonymous):

now you can ask me. I have used integrate by parts.

OpenStudy (anonymous):

sorry i have not done by linear approximation.